Consider two simple pendulums P and Q of equal length L and masses m1 and m2 respectively. Suppose the bobs of the pendulum are connected to each other by a weightless spring of force constant k such that in the equilibrium position, the spring is unstretched. Calculate the normal mode frequencies, and draw labelled diagrams corresponding to each mode.
1
Expert's answer
2020-02-28T10:09:33-0500

In short,


the first mode of oscillation


"\omega_1=\sqrt{\frac{g}{L}}"





the second mode of oscillation


"\omega_2=\sqrt{\frac{g}{L}+k\cdot \frac{m_1+m_2}{m_1\cdot m_2}}"